# Frequency-Aware Gaussian Splatting Decomposition

### v1

@@ Abstract
3D Gaussian Splatting (3D-GS) has revolutionized novel view synthesis with its efficient, explicit representation. However, it lacks frequency interpretability, making it difficult to separate low-frequency structures from fine details. We introduce a frequency-decomposed 3D-GS framework that groups 3D Gaussians that correspond to subbands in the Laplacian Pyrmaids of the input images.
Our approach enforces coherence within each subband (i.e., group of 3D Gaussians) through dedicated regularization, ensuring well-separated frequency components. We extend color values to both positive and negative ranges, allowing higher-frequency layers to add or subtract residual details. To stabilize optimization, we employ a progressive training scheme that refines details in a coarse-to-fine manner.
Beyond interpretability, this frequency-aware design unlocks a range of practical benefits. Explicit frequency separation enables advanced 3D editing and stylization, allowing precise manipulation of specific frequency bands. It also supports dynamic level-of-detail control for progressive rendering, streaming, foveated rendering and fast geometry interaction. Through extensive experiments, we demonstrate that our method provides improved control and flexibility for emerging applications in scene editing and interactive rendering. Our code will be made publicly available.
Figure 1: Frequency-Aware Decomposition for 3D Gaussian Splatting: We introduce a frequency-aware decomposition for 3D Gaussian Splatting (3D-GS). The decomposition assigns different Gaussians to different levels of the Laplacian Pyramid of the input images. This enables Level of Detail (LOD) rendering, Frequency-aware 3D editing, Foveated rendering, and 3D artistic filters.
@@ 1 Introduction
Novel view synthesis has undergone a remarkable evolution in recent years, propelled by the success of neural representation and optimization. First, Neural Radiance Fields (NeRFs) [10] and later 3D Gaussian Splatting (3D-GS) [5] have enabled the creation of high-fidelity scene reconstructions and photorealistic renderings from sparse input views.
Adding multi-resolution capabilities to 3D-GS representation can greatly help in tasks such as anti-aliasing, Level-Of-Details, or streaming, and considerable amount of work was done in this direction. Here, we take this idea one step further and merge 3D-GS with Laplacian Pyramids of images.
Specifically, we propose to group Gaussians based on the frequency sub-bands of the Laplacian pyramid of the input images. This allows us to create a structured hierarchy within 3D-GS, where each group of 3D Gaussians is used to render a specific sub-band across all input images. This design enables us to selectively render low-pass or high-pass components at will, offering a controllable frequency decomposition analogous to classical 2D image pyramids, but directly in 3D. Rendering only the first $k$ sub-bands yields a low-frequency rendering of the scene, while including all sub-bands produces a high-fidelity, full-frequency representation comparable to standard 3D-GS.
Beyond the fundamental advantage of greater interpretability, our frequency-aware 3D-GS representation opens the door to novel applications. Similar to level-of-detail systems, it naturally supports progressive rendering and streaming, allowing incremental refinement of a scene. Furthermore, it facilitates foveated rendering (rendering high-frequency detail only in a region of interest), frequency-aware 3D editing, scene stylization via consistent artistic filters that manipulate specific frequency ranges, and faster interaction with the scene’s geometry. We demonstrate that these enhancements go beyond simple scene reconstruction, highlighting the broad utility of a frequency-decomposed 3D Gaussian Splatting framework for both research and practical deployment. Fig. 1 demonstrates some of the mentioned applications.
In summary, our main contribution is a frequency-aware 3D Gaussian splatting decomposition, that enables explicit control over scene details across different frequency levels. This decomposition facilitates various applications, including progressive rendering, streaming, foveated rendering, geometry interaction, and advanced 3D editing.
@@ 2 Related Work
Figure 2: Architecture Overview: (Left) We group 3D Gaussians based on their contribution to different frequency levels and progressively accumulate them to construct a multi-scale representation. The accumulated contribution of the first $k$ Gaussian groups is designed to match a corresponding low-pass filtered version of the ground truth image, ensuring that each individual level aligns with the associated band in the Laplacian pyramid of the input images. (Middle) The training pipeline enforces this frequency-aware decomposition by supervising projections at different levels against progressively blurred ground truth images. (Right) We incorporate a frequency-domain loss, combining pixel-wise supervision with Discrete Fourier Transform (DFT) magnitude regularization to ensure that each frequency level contributes meaningfully to the final rendered image, preserving structural integrity while capturing high-frequency details.
@@ 2.1 Novel View Synthesis
Novel view synthesis aims to generate photorealistic images of a scene from previously unseen viewpoints, typically using only a sparse set of reference images as input. Early approaches in this field often relied on explicit scene geometry, such as depth-based image warping [13], to synthesize novel views. More recently, Neural Radiance Fields (NeRF) [10] have revolutionized the field by representing a 3D scene as a continuous function, parameterized by a multi-layer perceptron (MLP), that predicts color and density at any 3D coordinate and viewing direction. This enables high-quality view synthesis by querying the scene using volume rendering. However, despite their success, NeRF-based methods are computationally expensive, making both training and real-time rendering challenging. To address these limitations, a more recent approach, 3D Gaussian Splatting (3D-GS) [5], represents the scene as a set of explicit, oriented Gaussian primitives and employs a fast differentiable rasterizer to render novel views. This technique significantly reduces computational overhead while maintaining high-quality rendering results.
@@ 2.2 Frequency-Aware Representations
Many computer vision and graphics algorithms leverage frequency decomposition techniques to process and analyze images effectively. Classical methods such as the Laplacian pyramid [3] enable multi-scale image representations, facilitating detail manipulation and compression. The Image Fourier transform decomposes signals into frequency components, widely used in filtering and texture analysis, while wavelet transforms [7] provide localized frequency representations useful for image denoising and compression. Inspired by these techniques, recent neural representations incorporate frequency awareness to improve learning and reconstruction quality. [18] introduce Polynomial Neural Fields (PNFs) and demonstrate frequency-based manipulation of neural fields using Fourier PNFs. [21] propose progressive frequency training to accelerate the training and rendering of neural radiance fields.
Similarly to our approach, [20] utilize a Fourier-based loss, but apply it to guide Gaussian densification in a coarse-to-fine manner to mitigate over-reconstruction. In contrast, we use it to group Gaussians into distinct frequency subbands of the scene.
@@ 2.3 Level of Detail in 3D Gaussian Splatting
Similar to classical computer vision techniques, some 3D Gaussian Splatting (3D-GS) methods incorporate level-of-detail representations for various tasks. [17] introduce multi-scale Gaussians to improve anti-aliasing and enhance rendering efficiency at lower resolutions. [12] propose an octree-based data structure [8] to hierarchically organize Gaussians with anchors, ensuring consistent rendering speed across different resolutions and scales, particularly in large scenes. However, while these methods incorporate level-of-detail mechanisms, they lack explicit frequency-based regularization as in our approach, leading to missing structural information rather than just coarser details at lower levels of detail.
@@ 3.2 Gaussian Level Grouping
Progressive Level Introduction: Every $K$ training steps, we introduce a new level to the scene by duplicating all Gaussians in the scene, and assigning the next frequency level to the duplicated Gaussians.
Fig. 3 shows the decomposition of 3D Gaussians to different groups that correspond to different levels of the Laplacian pyramid. As can be seen, the base level requires a considerable amount of Gaussians, while the intermediate low levels require far less, except for the last two levels, that correspond to high frequencies.
@@ 3.3 Residual Color Gaussians
Level 1 Gaussians retain colors in the range $[0,1]$, consistent with the original 3D-GS paper, by using view-dependent spherical harmonic color generation; Given Gaussian $G_{i}$ and viewing direction $d$, the rendered color of the Gaussian is calculated by:
@@
Inspired by the Laplacian pyramid [3] in image processing, where residual levels can take both positive and negative values to represent higher-frequency sub-bands, we multiply the color range of all Gaussians at higher levels by 2 and shift down by 1, resulting in a range of $[-1,1]$:
@@ 3.4 Progressive Rendering of Frequency Bands
As stated above, at inference time, one can render all Gaussians in the scene at once to get the full resolution rendered image. Recall, in the Laplacian Pyramid, one can reconstruct a Gaussian Pyramid by up-sampling and adding each Laplacian level with its successor, resulting in images with progressively finer details. Similarly, to regulate the different frequency levels in our method during the training stage, we render each level of the Gaussian Pyramid separately by combining all Gaussian groups up to the corresponding frequency level.
@@
Specifically, for a pyramid with $L$ levels, rendering the $k$’th frequency level, where $1\leq k\leq L$, we render all Gaussians $G_{i}$ that satisfy $l_{i}\leq k$, results with the following sequence of rendered images:
@@ 3.5 Frequency Regulation and Losses Design
This way, we obtain the frequency spectrum of the GT image as it was subsampled $L-k$ times, but we keep it in it’s original resolution.
Figure 4: Progressive Rendering and Streaming: PSNR as a function of Frames Per Second (FPS), achieved by rendering all levels up to $L_{k}$. The model was trained at a resolution of $3115\times 2078$ pixels with six frequency levels, on the kitchen scene from the Mip-NeRF 360 dataset. The marker size radius is proportional to the accumulated number of Gaussians rendered across levels, illustrating the trade-off between fidelity and performance. Rendering more Gaussians increases PSNR but reduces FPS.
@@
Employing Eq. 7, we will define our frequency magnitude loss by applying the frequency discrepency term on all levels, except for the last one (as we do not desire to limit its frequency band):
@@
Given an image $I$ and a Gaussian kernel $\mathcal{G}$, the REDUCE operation is defined as a convolution with the kernel followed by subsampling every second row and column:
@@
To regulate the $k$-th frequency level, we apply Eq. 9 $L-k$ times to both $I_{k}$ and $\hat{I}_{k}$, yielding:
@@
where $d_{\text{IM}}$ combines a pixel-wise difference term with a structural similarity (SSIM) discrepancy, weighted by $\lambda_{\text{SSIM}}$. Utilizing Eq. 11, the spatial image loss is computed by summing the spatial image discrepancy terms over the rendered and ground truth (GT) levels:
@@ 4 Results
We evaluate our novel representation by showcasing several applications that are not feasible with the standard 3D-GS implementation, highlighting the versatility and practical advantages of our frequency decomposition beyond conventional reconstruction and view synthesis.
4.1 Datasets
Following the 3D-GS dataset configuration, we evaluated our model across all nine scenes of the Mip-NeRF 360 dataset [1], the Truck and Train scenes from the Tanks & Temples dataset [6], and the Playroom and Dr. Johnson scenes from the Deep Blending dataset [4]. For evaluation, we adopt the same protocol as 3D-GS, selecting every eighth image as part of the test set.
@@ 4.2 Implementation details
All scenes are trained for 30,000 steps, with Gaussian refinement enabled in our model for the entire training process. Optimization is performed using SplatFacto’s optimizers, and all experiments are conducted on a single NVIDIA RTX A5000 GPU. For both our model and the baseline SplatFacto model, we incorporate the view-dependent anti-aliased opacity term proposed in Mip-Splatting [19].

### v2

@@ Abstract
3D Gaussian Splatting (3D-GS) enables efficient novel view synthesis, but treats all frequencies uniformly, making it difficult to separate coarse structure from fine detail. Recent works have started to exploit frequency signals, but lack explicit frequency decomposition of the 3D representation itself. We propose a frequency-aware decomposition that organizes 3D Gaussians into groups corresponding to Laplacian-pyramid subbands of the input images. Each group is trained with spatial frequency regularization to confine it to its target frequency, while higher-frequency bands use signed residual colors to capture fine details that may be missed by lower-frequency reconstructions. A progressive coarse-to-fine training schedule stabilizes the decomposition.
Our method achieves state-of-the-art reconstruction quality and rendering speed among all LOD-capable methods. In addition to improved interpretability, our method enables dynamic level-of-detail rendering, progressive streaming, foveated rendering, promptable 3D focus, and artistic filtering. Our code will be made publicly available.
Figure 1: Frequency-Aware Decomposition for 3D Gaussian Splatting: We introduce a frequency-aware decomposition for 3D Gaussian Splatting (3D-GS). The decomposition assigns different Gaussians to different levels of the Laplacian Pyramid of the input images. This enables Level of Detail (LOD) rendering, Promptable 3D focus, Foveated rendering, and 3D artistic filters. Our method achieves the highest PSNR, SSIM, and rendering speed among all LOD-capable baselines.
11footnotetext: Equal contribution.
@@ 1 Introduction
3D Gaussian Splatting (3D-GS) [9] is an explicit, rasterization-based scene representation that delivers real-time, photorealistic view synthesis. But it lacks a notion of frequency: the millions of Gaussians are an unstructured pool, and beyond heuristics, there is no principled way to say which Gaussians carry low-frequency structure and which encode fine detail. Recent works have begun to exploit frequencies for improved reconstruction quality - for example, FreGS [36] uses progressive frequency regularization to address over-reconstruction issues during densification. However, these approaches lack explicit frequency decomposition of the 3D representation itself, making it difficult to perform tasks such as level-of-detail rendering, efficient streaming, and control over different frequency components. Our goal is precisely to introduce this missing structural organization, which enables the above applications and more, as demonstrated in Frequency-Aware Gaussian Splatting Decomposition - Level of Detail: render the scene at a chosen level of detail for progressive rendering and streaming. Promptable 3D Focus: keep selected objects sharp while blurring the rest based on a text prompt, in a 3D-consistent and interactive manner. Foveated Rendering: reduce detail in the peripheral area, increasing rendering speed. 3D Artistic Filters: apply different artistic styles directly to the 3D scene.
We realize this structural organization by explicitly partitioning the Gaussians into a small number of groups that correspond to specific subbands in the input images, leveraging the 2D frequency bands of the training images to define these groups. We implement this through a progressive, coarse-to-fine training scheme. First, we optimize a base level of Gaussians to explain the low-frequency content of the images; we then introduce a new set of Gaussians for the next band and jointly optimize both levels, while constraining the earlier level to remain low pass using a frequency regularizer. This forces the newly added Gaussians to complement the reconstruction by capturing the missing higher-frequency residuals. Higher bands are added in the same manner. To model residuals faithfully, Gaussians in non-base levels use signed residual colors (RGB values can be positive or negative), allowing them to add or subtract color w.r.t. the accumulated rendering. The end result is a frequency-aware hierarchy in which each group renders a specific Laplacian-like subband, and any prefix of the hierarchy produces a valid, lower-detail reconstruction.
Using 2D image frequencies as a proxy is a deliberate design choice: it provides a stable and well-studied frequency basis, avoids the complexity of defining frequency for anisotropic splats that project differently across views, and, when enforced consistently across all views, yields a coherent segmentation of the underlying 3D content.
Our decomposition method achieves state-of-the-art reconstruction metrics among level-of-detail and frequency-aware 3D-GS methods on the standard benchmarks (Mip-NeRF360, Tanks&Temples, and Deep Blending) while preserving real-time rendering.
In short, we reframe level-of-detail not as mere hierarchical sparsification, but as frequency segmentation of the representation itself. Our contributions are: (i) a frequency-aware decomposition of 3D-GS driven by image-space Laplacian bands, (ii) a progressive training strategy with signed residual colors and band-separating regularization, and (iii) comprehensive experiments demonstrating SOTA quality among LOD/frequency-aware methods together with diverse, frequency-driven applications.
@@ 2 Related Work
Figure 2: Frequency-Aware 3D Gaussian Splatting Overview: We decompose 3D Gaussians into frequency-specific groups to construct a multi-scale representation. (Left) Each Gaussian group targets a specific frequency band: Level 1 captures low-frequency structure, while Levels 2-3 add mid and high-frequency details through signed residual colors. (Middle) During training, each accumulated level (L1, L1+L2, L1+L2+L3) is supervised against appropriately blurred ground truth images, with red dashed lines showing gradient flow that constrains lower levels to remain frequency-limited while higher levels capture residual details. (Right) Our dual loss combines standard pixel loss with DFT magnitude loss in frequency space, ensuring each level contributes meaningful frequency content and preventing frequency leakage between bands.
@@ 2.1 Novel View Synthesis
Novel view synthesis aims to generate photorealistic images from previously unseen viewpoints, typically given only sparse reference views. Early approaches primarily relied on explicit geometry or depth-based warping, such as layered depth images [24], view interpolation [6], and light field modeling [13]. Later methods incorporated learned priors and convolutional neural networks to predict novel viewpoints directly from input images [38, 21]. A significant breakthrough came with Neural Radiance Fields (NeRF) [19], which represent scenes as continuous volumetric functions parameterized by multi-layer perceptrons, enabling high-quality view synthesis via volume rendering. NeRF has since inspired a series of high‑profile successors  [1, 20, 3] that extend the original framework. However, the implicit MLP representation common to these models still imposes a non‑trivial computational load. Addressing efficiency from a different angle, 3D Gaussian Splatting (3D‑GS) [9] replaces the network with explicit oriented Gaussians and a fast differentiable rasterizer, markedly reducing overhead while preserving high‑quality rendering.
@@ 2.2 Frequency-Aware Representations
Many classical vision and graphics methods rely on frequency decompositions, such as Laplacian pyramids [5], discrete wavelet transforms [16], and Fourier-domain filtering, enabling multiscale analysis. [33] introduce Polynomial Neural Fields (PNFs) and demonstrate frequency-based manipulation of neural fields using Fourier PNFs. Neural rendering methods integrate similar concepts using Fourier-feature embeddings [28], progressive-frequency training [39], frequency-aware NeRFs for enhancing fine detail [37] and wavelet-based NeRF architectures [31, 11].
Within 3DGS, recent methods have started to exploit frequency signals. Mip‑Splatting [34] enforces alias‑free rendering via analytic Gaussian filtering based on sampling rate limits. GES  [7] additionally considers frequency aware loss using DoG.  [35] links Gaussian scale and density to underlying frequency content, and  [26] extends this notion to dynamic scenes. Similarly to our approach, FreGS [36] utilizes a Fourier-based loss, but applies it to guide Gaussian densification in a coarse-to-fine manner to mitigate over-reconstruction. In contrast, we use it to group Gaussians into distinct frequency subbands of the scene.
@@ 2.3 Level of Detail in 3D Gaussian Splatting
Recent works have extended 3D Gaussian Splatting with level-of-detail mechanisms. Multi-Scale 3D-GS [32] introduces multiple Gaussian scales to reduce aliasing artifacts low resolutions, while OCTree-GS [22] organizes Gaussians within an octree structure [17] to accelerate rendering across multiple scales and viewpoints. Other methods, including  [10, 14, 15, 30], organize large-scale scenes using hierarchical data structures for efficient rendering, but focus primarily on scalability and massive scene handling rather than explicit frequency or detail separation.
LapisGS [27] addresses adaptive streaming by training Gaussian layers sequentially on progressively higher-resolution inputs, allowing opacity changes between layers to enforce a coarse-to-fine structure. FLoD [23] organizes Gaussians into discrete levels according to their spatial extent, enabling customized resolution control through sequential optimization. Both approaches freeze most of the parameters of previously optimized levels before introducing new ones, relying on heuristic measures: resolution or spatial extent, to define their level-of-detail separation.
In contrast, our method explicitly defines Gaussian groups based on the frequency content of input images, jointly training all frequency groups progressively rather than sequentially freezing lower levels. We employ dedicated frequency-domain regularization to enforce clear spectral boundaries between these groups. Additionally, higher-frequency Gaussians use signed residual colors to encode fine details efficiently by subtracting colors from coarser frequency levels, similar to a Laplacian pyramid, thereby reducing the total number of Gaussians required.
@@ 3.2 Gaussian Level Grouping
Progressive Level Introduction: Every $K$ training steps, we introduce a new level by duplicating all Gaussians and assigning them the next frequency level; almost all duplicates are removed during training (see Supplementary for Gaussian count dynamics).
Fig. 3 shows the decomposition of 3D Gaussians to different groups that correspond to different levels of the Laplacian pyramid. As can be seen, the base level requires a considerable amount of Gaussians, while the intermediate low levels require far less, except for the last two levels, which correspond to high frequencies.
@@ 3.3 Residual Color Gaussians
Level 1 Gaussians retain colors in the range $[0,1]$, consistent with the original 3D-GS paper, by using view-dependent spherical harmonic color generation; Given a Gaussian $G_{i}$ and viewing direction $d$, the rendered color of the Gaussian is computed as:
@@
Inspired by the Laplacian pyramid [5] in image processing, where residual levels can take both positive and negative values to represent higher-frequency sub-bands, we multiply the color range of all Gaussians at higher levels by 2 and shift down by 1, resulting in a range of [1,1][-1,1]:
@@ 3.4 Progressive Rendering of Frequency Bands
As stated above, at inference time, one can render all Gaussians in the scene at once to get the full resolution rendered image. Recall, in the image Laplacian Pyramid, one can reconstruct a Gaussian Pyramid by up-sampling and adding each Laplacian level with its successor, resulting in images with progressively finer details. Similarly, to regulate the different frequency levels in our method during the training stage, we render each level of the Gaussian Pyramid separately by combining all Gaussian groups up to the corresponding frequency level.
@@
Specifically, for a pyramid with $L$ levels, rendering the $k$-th frequency level, where $1\leq k\leq L$, we render all Gaussians $G_{i}$ that satisfy $l_{i}\leq k$, resulting in the following sequence of rendered images:
@@ 3.5 Frequency Regulation and Losses Design
This way, we obtain the frequency spectrum of the GT image as it was subsampled LkL-k times, but we keep it in its original resolution.
Method Mip-NeRF360 Tanks & Temples Deep Blending
[…]
Lapis-GS** 0.753 26.32 0.248 6252724 62.7 0.811 23.27 0.215 2864263 63.70 0.892 28.94 0.222 3874530 48.1
[…]
Table 1: Comparison between our method and competing methods across three datasets. Cell colors indicate first (red), second (orange), and third (yellow) place for each metric. Our method ranks in the top 3 across all metrics and datasets, achieving the best SSIM, PSNR, and FPS in every case. * Indicates results taken from the original paper, not re-evaluated locally. ** Lapis-GS was trained with $\lambda_{\text{SSIM}}=0.2$ following their suggestion to overcome GPU OOM (OOM on a 48GB Nvidia A6000).
@@
Employing Eq. 7, we will define our frequency magnitude loss by applying the frequency discrepancy term on all levels, except for the last one (as we do not desire to limit its frequency band):
@@
Given an image $I$ and a 2D Gaussian kernel $\mathcal{G}$, the REDUCE operation is defined as a convolution with the kernel followed by subsampling every second row and column:
@@
To regulate the $k$-th frequency level, we apply Eq. 9 LkL-k times to both $I_{k}$ and $\hat{I}_{k}$, yielding:
@@
where $d_{\text{IM}}$ combines a pixel-wise difference term with a structural similarity (SSIM) discrepancy $d_{\text{D-SSIM}}$, weighted by $\lambda_{\text{SSIM}}$. Utilizing Eq. 11, the spatial image loss is computed by summing the spatial image discrepancy terms over the rendered and ground truth (GT) levels:
@@ 4 Results
We evaluate our method across a comprehensive set of metrics and benchmarks, aiming to validate both reconstruction quality and real-time rendering efficiency. Specifically, we report SSIM, PSNR, LPIPS, Gaussian count, and FPS, and compare against prior 3D Gaussian Splatting methods that support level-of-detail (LOD) rendering. For fair comparison, we re-trained and evaluated baselines using the same evaluation protocol. In addition, we provide per-level progressive rendering analysis and demonstrate several applications uniquely enabled by our frequency-aware representation, including streaming, foveated rendering, and promptable 3D focus.
Data and Evaluations: Following the 3D-GS dataset configuration, we evaluated our model across all nine scenes of the Mip-NeRF 360 dataset [2], the Truck and Train scenes from the Tanks & Temples dataset [12], and the Playroom and Dr. Johnson scenes from the Deep Blending dataset [8]. For evaluation, we adopt the 3D-GS protocol, selecting every eighth image as part of the test set.
@@ 4.1 Implementation Details
All scenes are trained for 30,000 steps, with Gaussian refinement enabled in our model for the entire training process. Optimization is performed using SplatFacto’s optimizers, and all experiments are conducted on a single NVIDIA RTX A5000 GPU. For our model, we incorporate the view-dependent anti-aliased opacity term proposed in Mip-Splatting [34].
4.2 Comparisons with the State-of-the-Art
We compare our method against recent 3D Gaussian Splatting techniques, including FLOD [23], OCTree-GS [22], MSGS [32], FreGS [36], and Lapis-GS [27]. While most of these methods support explicit level-of-detail (LOD) rendering, FreGS is included for its use of frequency-aware modeling, even though it does not offer progressive rendering. We also include 3DGS [9] as a baseline. Wherever possible, we re-trained or re-evaluated competing methods locally using the same evaluation protocol.
We evaluated two variants of our method: Ours 3LOD, which uses three frequency levels, and Ours 5LOD, which uses five. Both are trained with the same budget of 30,000 steps. These configurations allow fine-grained trade-offs between quality and performance, and demonstrate the flexibility of our frequency decomposition.
As shown in Table 1, both our 3LOD and 5LOD configurations consistently rank among the top three methods across all metrics and datasets. Notably, Ours 3LOD achieves the highest SSIM, PSNR, and FPS in every dataset, establishing it as the best performing method in terms of both reconstruction quality and rendering speed. Our models maintain highly competitive LPIPS scores despite using significantly fewer Gaussians than dense baselines like 3DGS and FLOD.